# DCNv3: Towards Next Generation Deep Cross Network for Click-Through Rate Prediction

### v6

@@ 4.3.1. Efficiency Comparison (RQ2)
To verify the efficiency of DCNv3, we fix the optimal hyperparameters of the 25 baseline models and compare their parameter count (rounded to two decimal places) and runtime (averaged over five runs). The experimental results are shown in Figure 5. We can derive:
@@
  • Our proposed ECN and DCNv3 are the most parameter-efficient models in the DCN series, requiring only 0.78M and 1.37M parameters respectively to achieve SOTA performance. Meanwhile, regarding runtime, DCNv3 consistently outperforms strong baseline models such as FinalMLP, FINAL, DCNv2, and DCNv1. This demonstrates the time-efficient of DCNv3. Although DCNv3 requires an additional 20 seconds compared to ECN due to the employment of the Tri-BCE loss, it still remains comparable to DCNv2. Notably, the extra computational cost for the loss is only incurred during training and does not affect inference speed in practical applications. This further demonstrates the efficiency of ECN and DCNv3.
4.3.2. The Interpretability and noise filtering capability of DCNv3 (RQ3)
Interpretability is crucial in CTR prediction tasks (Li et al., 2020; Wang et al., 2023a, 2021b; Song et al., 2019) as it helps researchers understand predictions and increases confidence in the results. In this section, we investigate the dynamic Cross & Masked Vector and the static $\mathbf{W}_{l}$ to understand the model’s prediction process. The experimental and visualization results are shown in Figure 6 (Value denotes the Frobenius norm for each feature field), we can derive the following observations:
(a) $\mathbf{c}_{l}$ of LCN in DCNv3
(b) $\texttt{Mask}(\mathbf{c}_{l})$ of LCN in DCNv3
(c) $\mathbf{c}_{l}$ of ECN in DCNv3
(d) $\texttt{Mask}(\mathbf{c}_{l})$ of ECN in DCNv3
(e) $\mathbf{W}_{3}$ of LCN in DCNv3
(f) $\mathbf{W}_{3}$ of ECN in DCNv3
Figure 6. Visualization of field-wise value and sparsity for $\mathbf{c}_{l}$, $\texttt{Mask}(\mathbf{c}_{l})$, and $\mathbf{W}_{3}$ of DCNv3 on a random batch of ML-1M.
@@ 4.3.2. The Interpretability and noise filtering capability of DCNv3 (RQ3)
  • $\mathbf{c}_{l}$ and $\texttt{Mask}(\mathbf{c}_{l})$ exhibit complementary properties. When a feature field in $\mathbf{c}_{l}$ becomes important, its corresponding sparsity in $\texttt{Mask}(\mathbf{c}_{l})$ decreases, and vice versa (e.g., UserID, Occupation, Age fields). This demonstrates the effectiveness of our introduced Self-Mask operation, which further filters out noise by more aggressively assigning zero values to certain feature fields.
  • From Figure 6 (e, f), we observe that LCN and ECN capture different feature interaction information at the same layer. In LCN, $\mathbf{W}_{3}$ is used to compute the importance of 3-order features to generate 4-order feature interactions. In contrast, in ECN, $\mathbf{W}_{3}$ is used to compute the importance of $2^{2}$-order features to generate $2^{3}$-order feature interactions. Consequently, ECN shows reduced importance for UserID $\times$ Genres compared to LCN. This further proves the validity of DCNv3.
  • Overall, we observe that the importance of high-order feature interactions is lower than that of lower-order feature interactions, which is similarly reported in some works (Liu et al., 2020; Wang et al., 2023a; Zhu et al., 2023a). For example, in Figure 6, (f) has fewer bright red blocks compared to (e), and the blue blocks in (a) gradually darken as the number of layers increases, the situation is similar in (b).
@@ 4.3.3. Ablation Study (RQ4)
To investigate the impact of each component of DCNv3 on its performance, we conduct experiments on several variants of DCNv3:
  • w/o TB: DCNv3 with BCE instead of the Tri-BCE.
@@
The results of the ablation experiments are shown in Table 5. It is observed that both ECN and LCN exhibit some performance loss compared to DCNv3, which demonstrates the necessity of capturing both high-order and low-order feature interactions. LCN consistently underperforms compared to ECN, demonstrating the superiority of the exponential feature interaction method. Meanwhile, the variant w/o TB also leads to a certain degree of performance decline, particularly noticeable on KKBox. LayerNorm aims to ensure that the Self-Mask maintains a masking rate of around 0.5, so its removal results in destabilizing the masking rate, leading to some performance loss. This demonstrates the necessity and effectiveness of each component within DCNv3.
Table 5. Ablation study of DCNv3.
  Model Criteo ML-1M KKBox
@@ 4.3.4. Influence of Network Depths (RQ4)
From Figure 7, we observe that on the Criteo dataset, the model achieves optimal performance at a depth of 4 layers, indicating that ECN captures up to $2^{4}$-order feature interactions. On the KDD12 dataset, ECN achieves optimal performance at a depth of 6 layers, meaning it captures $2^{6}$-order feature interactions. In contrast, achieving the same order of feature interactions in the linearly growing CrossNetv2 requires $12^{4}-$ 1 and $12^{6}-$ 1 layers, respectively. Considering the huge computational resources required, this is impractical, whereas ECN easily accomplishes this with its exponentially growing feature interaction mechanism. This further demonstrates the effectiveness of ECN.
@@ 5. Conclusion
This paper introduced the next generation deep cross network, called DCNv3, which uses sub-networks LCN and ECN to capture both low-order and high-order feature interactions without relying on the less interpretable DNN. LCN uses a linearly growing interaction method for low-order interactions, while ECN employs an exponentially increasing method for high-order interactions. The Self-Mask filters interaction noise and further improves DCNv3’s computational efficiency. Tri-BCE helped the two sub-networks in DCNv3 obtain more suitable supervision signals for themselves. Comprehensive experiments on six datasets demonstrated the effectiveness, efficiency, and interpretability of DCNv3. Additionally, our proposed models achieved 1st rankings in multiple CTR benchmarks using only explicit feature interactions, breaking the convention that traditional CTR models integrate implicit feature interactions to improve performance.

### v7

@@ 4.3.1. Efficiency Comparison (RQ2)
To verify the efficiency of FCN, we fix the optimal hyperparameters of the 25 baseline models and compare their parameter count (rounded to two decimal places) and runtime (averaged over five runs). The experimental results are shown in Figure 5. We can derive:
@@
  • Our proposed ECN and FCN are the most parameter-efficient models in the DCN series, requiring only 0.78M and 1.37M parameters respectively to achieve SOTA performance. Meanwhile, regarding runtime, ECN consistently outperforms strong baseline models such as FinalMLP, FINAL, DCNv2, and DCNv1. This demonstrates the time-efficient of ECN. Although FCN requires an additional 20 seconds compared to ECN due to the employment of the Tri-BCE loss, it still remains comparable to DCNv2. Notably, the extra computational cost for the loss is only incurred during training and does not affect inference speed. This further demonstrates the efficiency of ECN and FCN.
4.3.2. The Interpretability and noise filtering capability of FCN (RQ3)
Interpretability is crucial in CTR prediction tasks (Li et al., 2020; Wang et al., 2023a, 2021b; Song et al., 2019) as it helps researchers understand predictions and increases confidence in the results. In this section, we investigate the dynamic Cross & Masked Vector and the static $\mathbf{W}_{l}$ to understand the model’s prediction process. Formally, both the Cross & Masked Vector and $\mathbf{W}_{l}$ can be represented in a field-wise manner:
(10) $\mathbf{c}_{l}\ ||\ \texttt{Mask}(\mathbf{c}_{l})$ $=\left[c_{1},c_{2},\dots,c_{\scriptscriptstyle\frac{fd}{2}},m_{1},m_{2},\dots,m_{\scriptscriptstyle\frac{fd}{2}}\right],$
$=\left[C_{1},C_{2},\dots,C_{f},M_{1},M_{2},\dots,M_{f}\right],$
(11) $\mathbf{W}_{l}=\left[\begin{array}[]{ccc}w_{1,1}&\cdots&w_{1,fd}\\ \vdots&\ddots&\vdots\\ w_{\frac{fd}{2},1}&\cdots&w_{\frac{fd}{2},fd}\end{array}\right]=\left[\begin{array}[]{ccc}W_{1,1}&\cdots&W_{1,f}\\ \vdots&\ddots&\vdots\\ W_{f,1}&\cdots&W_{f,f}\end{array}\right],$
where $c_{i}=\sum_{j=1}^{D}w_{i,j}x_{j}\in\mathbb{R}^{1}$ represent the $i$-th elements in $\mathbf{c}_{l}$, $x_{j}\in\mathbf{x}$, and $C_{i}\in\mathbb{R}^{\frac{d}{2}}$ represents all the features to be crossed for the $i$-th feature field. From the perspective of attention mechanisms, the larger $\sum_{j=1}^{D}w_{i,j}$ is the higher the attention score, thus $C_{i}$ reflects the strength of interaction required for each feature field through soft selection. Similarly, $M_{i}\in\mathbb{R}^{\frac{d}{2}}$ represents the features to be crossed for the $i$-th feature field after hard selection via Self-Mask, with the sparsity of $M_{i}$ indicating the proportion of noise that needs to be filtered for each feature field. Furthermore, $w_{i,j}\in\mathbb{R}^{1}$ denotes the element in the $i$-th row and $j$-th column of $\mathbf{W}_{l}$. Each feature field is associated with a corresponding field-wise cross matrix $W_{i,j}\in\mathbb{R}^{\frac{d}{2}\times d}$, which represents the importance of interaction between the $i$-th and $j$-th feature fields. Typically, CTR researchers believe that as the number of cross layers increases, the corresponding cross matrices can further reveal the inherent relationships and importance among multiple feature fields (Wang et al., 2021b, 2023a).
The experimental and visualization results are shown in Figure 6 (Value denotes the Frobenius norm for each feature field), we can derive the following observations:
(a) $\mathbf{c}_{l}$ of LCN in FCN
(b) $\texttt{Mask}(\mathbf{c}_{l})$ of LCN in FCN
(c) $\mathbf{c}_{l}$ of ECN in FCN
(d) $\texttt{Mask}(\mathbf{c}_{l})$ of ECN in FCN
(e) $\mathbf{W}_{3}$ of LCN in FCN
(f) $\mathbf{W}_{3}$ of ECN in FCN
Figure 6. Visualization of field-wise value and sparsity for $\mathbf{c}_{l}$, $\texttt{Mask}(\mathbf{c}_{l})$, and $\mathbf{W}_{3}$ of FCN on a random batch of ML-1M.
@@ 4.3.2. The Interpretability and noise filtering capability of FCN (RQ3)
  • When a feature field in $\mathbf{c}_{l}$ becomes important, its corresponding sparsity in $\texttt{Mask}(\mathbf{c}_{l})$ decreases, and vice versa (e.g., UserID, Occupation, Age fields). This demonstrates the effectiveness of our introduced Self-Mask operation, which further filters out noise by more aggressively assigning zero values to certain feature fields. This indicates that $\mathbf{c}_{l}$ and $\texttt{Mask}(\mathbf{c}_{l})$ are complementary and reflect the same feature interaction patterns.
  • From Figure 6 (e, f), we observe that LCN and ECN capture different feature interaction information at the same layer. In LCN, $\mathbf{W}_{3}$ is used to compute the importance of 3-order features to generate 4-order feature interactions. In contrast, in ECN, $\mathbf{W}_{3}$ is used to compute the importance of $2^{2}$-order features to generate $2^{3}$-order feature interactions. From a local perspective, taking the UserID $\times$ Genres interaction pair as an example, we find that its contribution to the final prediction results is higher in LCN but lower in ECN. From a global perspective, the importance of interaction pairs in LCN is more distinct, while in ECN, the importance is more evenly distributed. This difference suggests that LCN excels at capturing the key features interaction of specific feature pairs, highlighting strong local associations, whereas ECN tends to balance all interaction pairs, enabling it to capture potential relationships between features more comprehensively and improve generalization ability. This complementarity plays an indispensable role in the final prediction results of FCN.
  • In Figure 6, subplot (f) contains fewer bright red blocks compared to subplot (e), while the blue blocks in subplot (a) gradually darken as the number of layers increases, with a similar trend observed in subplot (b). This indicates that the importance of high-order feature interactions is lower than that of lower-order feature interactions, a phenomenon also reported in previous works (Liu et al., 2020; Wang et al., 2023a; Zhu et al., 2023a). Overall, we can understand the learning process of FCN from the following three aspects: the interaction strength between feature fields (subplot (a) and (c)), the sparsity of features to be interacted (subplot (b) and (d)), and the importance of feature pair interactions (subplot (e) and (f)). This not only helps analyze the contribution of features and the noise involved during interactions but also aids in identifying which feature pairs have a significant impact on the model’s predictions. Once these key features are identified, we can further explain their roles and even intervene in related features to enhance the probability of user clicks.
@@ 4.3.3. Ablation Study (RQ4)
To investigate the impact of each component of FCN on its performance, we conduct experiments on several variants of FCN:
  • w/o TB: FCN with BCE instead of the Tri-BCE.
@@
The results of the ablation experiments are shown in Table 5. It is observed that both ECN and LCN exhibit some performance loss compared to FCN, which demonstrates the necessity of capturing both high-order and low-order feature interactions. LCN consistently underperforms compared to ECN, demonstrating the superiority of the exponential feature interaction method. Meanwhile, the variant w/o TB also leads to a certain degree of performance decline, particularly noticeable on KKBox. LayerNorm aims to ensure that the Self-Mask maintains a masking rate of around 0.5, so its removal results in destabilizing the masking rate, leading to some performance loss. This demonstrates the necessity and effectiveness of each component within FCN. Moreover, we find that simply combining multiple sub-networks does not always lead to performance improvements. For example, in KKBox and iPinYou datasets, the variant w/o TB performs worse than a single ECN, which has also been reported in some related works (Li et al., 2024b; Zhu et al., 2023b). This further demonstrates the effectiveness of our proposed Tri-BCE.
Table 5. Ablation study of FCN.
  Model Criteo ML-1M KKBox iPinYou
@@ 4.3.4. Influence of Network Depths (RQ4)
From Figure 7, we observe that on the Criteo dataset, the model achieves optimal performance at a depth of 4 layers, indicating that ECN captures up to $2^{4}$-order feature interactions. On the KDD12 dataset, ECN achieves optimal performance at a depth of 6 layers, meaning it captures $2^{6}$-order feature interactions. In contrast, achieving the same order of feature interactions in the linearly growing CrossNetv2 requires $12^{4}-$ 1 and $12^{6}-$ 1 layers, respectively. Considering the huge computational resources required, this is impractical, whereas ECN easily accomplishes this with its exponentially growing feature interaction mechanism. This further demonstrates the effectiveness of ECN. Besides, the implementation of FCN uses dropout (Srivastava et al., 2014), which is a common trick to prevent overfitting in many CTR model (Song et al., 2019; Zhu et al., 2021). We recommend setting the dropout rate to 0.1 regardless of the network’s depth, as it is a good choice. A dropout rate that is too high may cause the model to converge more slowly.
@@ 5. Conclusion
This paper introduced the next generation deep cross network, called FCN, which uses sub-networks LCN and ECN to capture both low-order and high-order feature interactions without relying on the less interpretable DNN. LCN uses a linearly growing interaction method for low-order interactions, while ECN employs an exponentially increasing method for high-order interactions. The Self-Mask filters interaction noise and further improves FCN’s computational efficiency. Tri-BCE helped the two sub-networks in FCN obtain more suitable supervision signals for themselves. Comprehensive experiments on six datasets demonstrated the effectiveness, efficiency, and interpretability of FCN. Additionally, our proposed models achieved 1st rankings in multiple CTR benchmarks using only explicit feature interactions, breaking the convention that traditional CTR models integrate implicit feature interactions to improve performance.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.